what is -4 square root 27 + square root 12
Answer by fcabanski(1391)

Check my profile and contact me for one on one online tutoring.


-4sqrt(27) + sqrt(12)


Remember sqrt(xy)=sqrt(x)*sqrt(y). Break up (factor) 27 and 12 into perfect square factors, so you can bring them out of the sqrt().


-4 sqrt(3)*sqrt(9) + sqrt(3)*sqrt(4) =


sqrt(9)=3 and sqrt(4)=2 so:


-4*3*sqrt(3) +2*sqrt(3) = -12sqrt(3) + 2sqrt(3) = -10sqrt(3) = approx. -17.32